# MGAug: Multimodal Geometric Augmentation in Latent Spaces of Image Deformations

### 20 Dec 2023 v1

@@ Abstract
Geometric transformations have been widely used to augment the size of training images. Existing methods often assume a unimodal distribution of the underlying transformations between images, which limits their power when data with multimodal distributions occur. In this paper, we propose a novel model, Multimodal Geometric Augmentation (MGAug), that for the first time generates augmenting transformations in a multimodal latent space of geometric deformations. To achieve this, we first develop a deep network that embeds the learning of latent geometric spaces of diffeomorphic transformations (a.k.a. diffeomorphisms) in a variational autoencoder (VAE). A mixture of multivariate Gaussians is formulated in the tangent space of diffeomorphisms and serves as a prior to approximate the hidden distribution of image transformations. We then augment the original training dataset by deforming images using randomly sampled transformations from the learned multimodal latent space of VAE. To validate the efficiency of our model, we jointly learn the augmentation strategy with two distinct domain-specific tasks: multi-class classification on 2D synthetic datasets and segmentation on real 3D brain magnetic resonance images (MRIs). We also compare MGAug with state-of-the-art transformation-based image augmentation algorithms. Experimental results show that our proposed approach outperforms all baselines by significantly improved prediction accuracy. Our code is publicly available at https://github.com/tonmoy-hossain/MGAug.
Introduction
@@
Recent advances in deep learning (DL) have led to remarkable progress in image analysis tasks, including but not limited to image classification (Krizhevsky, Sutskever, and Hinton 2017; Wang and Zhang 2022), segmentation (Hossain et al. 2019; Chen et al. 2021), object detection (Kim and Pavlovic 2016; Pang et al. 2019), and 3D reconstruction (Jayakumar, Hossain, and Zhang 2023). The literature has shown that deep neural networks (DNNs) are most powerful when trained on vast amounts of manually labeled data. Unfortunately, many image applications have relatively small annotated datasets available, due to privacy concerns, regulations, costs, and incidents of disease. The DNNs, in such cases, tend to overfit due to their poor generalizations to large data variances. Data augmentation (DA) is a simple, yet efficient and commonly used strategy to overcome this issue by synthesizing more training examples to artificially increase the dataset (Cubuk et al. 2019; Zhao et al. 2019; Zoph et al. 2020; Shen et al. 2020). It is often employed as a preprocessing step applied to the training data to expose the DNNs to various types of approximated invariance; hence reducing the problem of overfitting. In the image domain, common augmentation strategies include adding variations to image intensities (Gao et al. 2021), or deforming the image by randomly generated geometric transformations (i.e., translation, rotation, or scaling) (Moreno-Barea et al. 2018). This paper focuses on DA, particularly in geometric spaces.
@@
In practice, DA is usually a manual process where a human specifies a set of affine or deformable transformations believed to make DL models invariant (Hendrycks et al. 2019; Yun et al. 2019). However, such methods can pose a high risk of producing unrealistic or implausible variations of the image. For example, rotating an image 180 degrees makes sense for natural images, but produces ambiguities on digital numbers (i.e., the digits ‘6’ and ‘9’). In medical domain, randomly generated deformations may transform images from a healthy to a diseased group. To alleviate this issue, researchers have developed a group of automatic DA methods that learn an optimal augmenting policy from the data itself (Cubuk et al. 2019; Lim et al. 2019; Zoph et al. 2020; Araslanov and Roth 2021; Muralikrishnan et al. 2022). One approach is to learn a generator via generative adversarial networks (GANs) to synthesize data from scratch (Mirza and Osindero 2014; Chen et al. 2022). However, these models are difficult to train due to the high complexity of largely unknown parameter spaces (Jaderberg et al. 2015). Despite showing encouraging results on relatively easy datasets (e.g., $28^{2}$ MNIST), they do not appear to be applicable to more complex datasets with large structural and geometric variances. Another common approach is to learn a probabilistic distribution of a set of deformable transformations derived from the training images (Hauberg et al. 2016; Krebs et al. 2019; Dalca et al. 2019; Zhao et al. 2019; Olut et al. 2020; Shen et al. 2020). Most importantly, these methods assume a unimodal distribution of the underlying transformations between images. They do not provide an appropriate way to learn the full space of geometric transformations among the training data. This greatly limits their power when multimodal data (e.g., time-series/longitudinal images, or images of multiple sub-populations) occurs.
@@
In this paper, we present a novel model, MGAug, that for the first time learns augmenting transformations in a multimodal latent space of geometric deformations. In contrast to existing approaches limited to data with unimodal distributions, our MGAug features the main advantages of
  1. (i)
@@
We demonstrate the effectiveness of our proposed model in two distinct domain-specific tasks: multi-class classification and segmentation on both 2D synthetic data and real 3D brain magnetic resonance images (MRIs), respectively. We also compare MGAug with the state-of-the-art transformation-based augmentation algorithms (Olut et al. 2020; Dalca et al. 2019). Experimental results show that our method outperforms all baseline algorithms by significantly improved prediction accuracy.
Background: Geometric Deformations via Template-based Image Registration
We first briefly review the concept of template-based deformable image registration (Joshi et al. 2004; Zhang, Singh, and Fletcher 2013), which is widely used to estimate deformable geometric deformations from groupwise images. In this work, we adopt the algorithm of large deformation diffeomorphic metric mapping (LDDMM), which encourages the smoothness of deformation fields and preserves the topological structures of images (Beg et al. 2005).
@@
where $\sigma^{2}$ is a noise variance and $\circ$ denotes an interpolation operator that deforms image $I$ with an estimated transformation $\phi_{n}$. The $\text{Dist}[\cdot,\cdot]$ is a distance function that measures the dissimilarity between images, i.e., sum-of-squared differences (Beg et al. 2005), normalized cross correlation (Avants et al. 2008), and mutual information (Wells III et al. 1996). The ${\rm Reg}[\cdot]$ is a regularizer that guarantees the smoothness of transformations, i.e.,
@@
The minimum of Eq. (2) is uniquely determined by solving a Euler-Poincaré differential equation (EPDiff) (Arnold 1966; Miller, Trouvé, and Younes 2006) with a given initial condition of velocity fields, noted as $v_{n}(0)$,
@@
where the operator $D$ denotes a Jacobian matrix, $\operatorname{div}$ is the divergence, and $\cdot$ represents an element-wise matrix multiplication. This is known as the geodesic shooting algorithm (Vialard et al. 2012), which nicely proves that the deformation-based shape descriptor, $\phi_{n}$, can be fully characterized by an initial velocity field, $v_{n}(0)$.
@@
Our Method: MGAug
This section presents a novel algorithm, MGAug, that learns a multimodal latent distribution of geometric transformations represented by the initial velocity fields $v_{0}$. We first introduce a generative model in latent geometric spaces for DA based on VAE, followed by an approximated variational inference. The augmented data is then used to train various image analysis tasks, i.e., image classification or segmentation. Similar to (Zhao et al. 2019; Chen et al. 2022), we employ a joint learning scheme to optimize the MGAug and IAT network parameters. To simplify the notation, we omit the time index of $v_{0}$ in the following sections.
Generative Model in Latent Multimodal Geometric Spaces
Given a group of $N$ input images $\mathbf{J}$ with a template $I$, we assume the latent geometric transformations $\mathbf{x}$ are generated from a number of $C$ clusters. The distribution of $\mathbf{x}$ can be formulated as a Gaussian mixture with means $\boldsymbol{\mu}$ and variances $\mathbf{\Sigma}$, specified by a VAE encoder parameterized by $\beta$, i.e., $p_{\theta}(\mathbf{x}|{\boldsymbol{\epsilon}})$, where $\boldsymbol{\epsilon}$ is a random number generated from $\mathcal{N}(0,\mathbb{I})$. That is to say, the VAE encoder outputs a set of $C$ means $\boldsymbol{\mu}$ and $C$ variances $\mathbf{\Sigma}$. A one-hot vector $\mathbf{z}$ is sampled from the mixing probability $\mathbf{\boldsymbol{\pi}}\in[0,1]$ with $\sum_{c=1}^{C}\mathbf{\boldsymbol{\pi}}^{c}=1$, which chooses one component from the Gaussian mixture. In this paper, we set the parameter $\mathbf{\boldsymbol{\pi}}=C^{-1}$ to make the variable $\mathbf{z}$ uniformly distributed, which is similar to (Dilokthanakul et al. 2016). The initial velocity field $\mathbf{v}$ is generated from a VAE decoder parameterized by $\theta$ and the continuous latent variable $\mathbf{x}$. The entire process to generate the observed image data $\mathbf{J}$ from a set of latent variables $\{\mathbf{v},\mathbf{x},\mathbf{\boldsymbol{\epsilon}},\mathbf{z}\}$ is:
@@
Similar to (Zhang and Fletcher 2013; Zhang, Shao, and Fletcher 2015), we define the covariance matrix $\boldsymbol{\Sigma}$ as an inverse operator of the differential operator, $L$, which is used in the registration regularity in Eq. (2).
@@
Variational Lower Bound
@@
Figure 2: An overview of our model MGAug for image analysis tasks.
Using previously defined Eq. (10) and Eq. (11), the posterior $q(\mathbf{v}_{i}|\mathbf{J}_{i},I,\boldsymbol{\lambda})\propto p(\mathbf{J}|\mathbf{v},I,\mathbf{\boldsymbol{\lambda}})\cdot p(\mathbf{v}|\mathbf{x})$ is formulated by Baye’s rule. Following (Jiang et al. 2016), we can derive the posterior $p_{\beta}(\mathbf{z}|\mathbf{x},\mathbf{\boldsymbol{\epsilon}})$ as
@@
$-\mathbb{E}_{q(\mathbf{x}|\mathbf{J})\,q(\boldsymbol{\epsilon}|\mathbf{J})}[KL(p_{\beta}(\mathbf{z}|\mathbf{x},\boldsymbol{\epsilon})\,||\,p(\mathbf{z}))]$ (14)
Joint Learning of MAGug With Image Analysis
We demonstrate the effectiveness of MGAug in image analysis tasks (IAT) by developing a joint learning framework. It has been shown in (Zhao et al. 2019; Chen et al. 2022) that these two tasks can be mutually beneficial. The MGAug effectively synthesizes image samples to augment the training dataset, while the IAT can provide extra information on rectifying the learned distributions of latent geometric spaces. An overview of our proposed network architecture is shown in Fig. 2.
@@
MGAug is mainly designed based on the previously introduced inference model, where the approximate posteriors are simulated by neural networks. Analogous to (Kingma and Welling 2013), we let the encoder of MGAug, parameterized by $\beta$, approximate the posterior as a mixture of multivariate Gaussians. In the latent space $\mathbf{z}$, we use non-linear hidden layers to learn the parameters (i.e., means and covariances) of the mixture Gaussians. The decoder, parameterized by $\theta$, outputs augmented velocity fields $\mathbf{v}_{n}$, followed by generating their associated deformations $\phi_{n}$ (via Eq. (3)) and augmented deformed images. In this paper, we adopt UNet architecture (Ronneberger, Fischer, and Brox 2015) as the network backbone for MGAug. However, other networks such as UNet++ (Zhou et al. 2018) and TransUNet (Chen et al. 2021) can be easily applied.
@@
$+\frac{1}{2}(L\mathbf{v}(\beta,\theta),\mathbf{v}(\beta,\theta))+\text{reg}(\beta,\theta),$
@@
For segmentation, we employ a loss function as a combination of Sørensen--Dice coefficient (Dice 1945) and cross-entropy (CE). Let $Q$ denote the number of segmentation labels, $S_{q}$ denotes the augmented ground truth, and $R_{q}$ denotes the prediction with the associated probabilistic map $R^{p}_{q}$ for $q$-th label. We formulate the segmentation loss as
@@
The total loss of our joint learning framework of MGAug and IAT is $\ell=\ell_{\text{MGAug}}+\ell_{\text{IAT}}$. We employ an alternative optimization scheme (Nocedal and Wright 1999) to minimize the loss by alternating between the training of the MGAug network and the IAT network without sharing weights, making it an end-to-end network architecture (Zhao et al. 2019; Chen et al. 2022).
Experimental Evaluation
@@
Datasets
2D silhouette shapes. We choose Kimia216 (Sebastian, Klein, and Kimia 2004), which is a relatively smaller dataset with seven classes of 2D silhouette shape data (bone, bottle, children, face, fish, heart, and key). Each shape class includes approximately $20$ binary images, with the size of $28^{2}$. The total number of our training images for multi-classification is about $140$.
2D handwritten digits. We use $70000$ handwritten digits of $10$ classes (00-$9$) from the MNIST repository (LeCun 1998). Each class includes approximately $7000$ images with the size of $28^{2}$.
3D brain MRIs. We include $803$ subjects of public T1-weighted longitudinal brain MRIs with manually delineated anatomical structures from Open Access Series of Imaging Studies (OASIS) (LaMontagne et al. 2019), Alzheimer’s Disease Neuroimaging Initiative (ADNI) (Jack Jr et al. 2008), Autism Brain Imaging Data Exchange (ABIDE) (Di Martino et al. 2014), and LONI Probabilistic Brain Atlas Individual Subject Data (LPBA40) (Shattuck et al. 2008). Due to the difficulty of preserving the diffeomorphic property across individual subjects with large age variations, we carefully evaluate images from subjects aged from $60$ to $90$. All MRIs were pre-processed as $256\times 256\times 256$, $1.25mm^{3}$ isotropic voxels, underwent skull-stripping, intensity normalization, bias field correction and pre-alignment with affine transformations.
Experiments and Implementation Details
We validate the proposed model, MGAug, on both image classification and segmentation tasks. As MGAug falls into the category of transformation-based DA schemes, we mainly focuses on comparing its performance with similar baselines: randomly generated affine transformations (ATA, a commonly used augmentation strategy), and two state-of-the-art unimodal generative models based on diffeomorphic registrations (AdvEigAug or AEA, an augmentation model based on the statistics of a learned latent space of velocity fields (Olut et al. 2020); and UDA, a unimodal deformation-based augmentation using random stationary velocity fields (Dalca et al. 2019)).
Evaluation of classification on small datasets. We first validate our model and other baselines on a relatively small dataset of 7-class 2D silhouette shapes. We adopt six classification network backbones, including a three-layer CNN, ResNet18 (He et al. 2016), VGGNet16 (Simonyan and Zisserman 2014), SqueezeNet (Iandola et al. 2016), DenseNet121 (Huang et al. 2017), and Vision Transformer (ViT) (Dosovitskiy et al. 2020).

### 25 Jan 2024 v2

@@ Abstract
Geometric transformations have been widely used to augment the size of training images. Existing methods often assume a unimodal distribution of the underlying transformations between images, which limits their power when data with multimodal distributions occur. In this paper, we propose a novel model, Multimodal Geometric Augmentation (MGAug), that for the first time generates augmenting transformations in a multimodal latent space of geometric deformations. To achieve this, we first develop a deep network that embeds the learning of latent geometric spaces of diffeomorphic transformations (a.k.a. diffeomorphisms) in a variational autoencoder (VAE). A mixture of multivariate Gaussians is formulated in the tangent space of diffeomorphisms and serves as a prior to approximate the hidden distribution of image transformations. We then augment the original training dataset by deforming images using randomly sampled transformations from the learned multimodal latent space of VAE. To validate the efficiency of our model, we jointly learn the augmentation strategy with two distinct domain-specific tasks: multi-class classification on 2D synthetic dataset and segmentation on real 3D brain magnetic resonance images (MRIs). We also compare MGAug with state-of-the-art transformation-based image augmentation algorithms. Experimental results show that our proposed approach outperforms all baselines by significantly improved prediction accuracy. Our code will be publicly available on GitHub.
I Introduction
@@
Recent advances in deep learning (DL) have led to remarkable progress in image analysis tasks, including but not limited to image classification [43, 38], segmentation [9, 47, 65], object detection [31, 49], and annotation [58, 10, 67]. The literature has shown that deep neural networks (DNNs) are most powerful when trained on vast amounts of manually labeled data. Unfortunately, many image applications have relatively small annotated datasets available, due to privacy concerns, regulations, costs, and incidents of disease. The DNNs, in such cases, tend to overfit due to their poor generalizations to large data variances. Data augmentation (DA) is a simple, yet efficient and commonly used strategy to overcome this issue by synthesizing more training examples to artificially increase the dataset [55, 11, 65, 68, 53]. It is often employed as a preprocessing step applied to the training data to expose the DNNs to various types of approximated invariance; hence reducing the problem of overfitting. In the image domain, common augmentation strategies include adding variations to the image intensities [23, 18], or deforming the image by randomly generated geometric transformations (i.e., translation, rotation, or scaling) [42]. This paper focuses on DA, particularly in geometric spaces.
@@
In practice, DA is usually a manual process where a human specifies a set of affine or deformable transformations believed to make DL models invariant [22, 61]. However, such methods can pose a high risk of producing unrealistic or implausible variations of the image. For example, rotating an image 180 degrees makes sense for natural images, but produces ambiguities on digital numbers (i.e., the digits ‘6’ and ‘9’). In medical domain, randomly generated deformations may transform images from a healthy to a diseased group. To alleviate this issue, researchers have developed a group of automatic DA methods that learn an optimal augmenting policy from the data itself [11, 36, 60, 68, 2, 45]. One approach is to learn a generator via generative adversarial networks (GANs) to synthesize data from scratch [40, 1, 37]. However, these models are difficult to train due to the high complexity of largely unknown parameter spaces [28]. Despite showing encouraging results on relatively easy datasets (e.g., $28^{2}$ MNIST), they do not appear to be applicable to more complex datasets with large structural and geometric variances. Another common approach is to learn a probabilistic distribution of a set of deformable transformations derived from the training images [20, 33, 13, 65, 48, 53]. Most importantly, these methods assume a unimodal distribution of the underlying transformations between images. They do not provide an appropriate way to learn the full space of geometric transformations among the training data. This greatly limits their power when multimodal data (e.g., time-series/longitudinal images, or images of multiple sub-populations) occurs.
@@
In this paper, we present a novel model, MGAug, that for the first time learns augmenting transformations in a multimodal latent space of geometric deformations. In contrast to existing approaches limited to data with unimodal distributions, our MGAug features main advantages of
  1. 1.
@@
We demonstrate the effectiveness of our proposed model in two distinct domain-specific tasks: multi-class classification and segmentation on both 2D synthetic data and real 3D brain magnetic resonance images (MRIs), respectively. We also compare MGAug with the state-of-the-art transformation-based augmentation algorithms [48, 13]. Experimental results show that our method outperforms all baseline algorithms by significantly improved prediction accuracy.
II Background: Geometric Deformations via Template-based Image Registration
We first briefly review the concept of template-based deformable image registration [30, 64], which is widely used to estimate deformable geometric deformations from groupwise images. In this work, we adopt the algorithm of large deformation diffeomorphic metric mapping (LDDMM), which encourages the smoothness of deformation fields and preserves the topological structures of images [5].
@@ II Background: Geometric Deformations via Template-based Image Registration
where $\sigma^{2}$ is a noise variance and $\circ$ denotes an interpolation operator that deforms image $I$ with an estimated transformation $\phi_{n}$. The $\text{Dist}[\cdot,\cdot]$ is a distance function that measures the dissimilarity between images, i.e., sum-of-squared differences [5], normalized cross correlation [4], and mutual information [59]. The ${\rm Reg}[\cdot]$ is a regularizer that guarantees the smoothness of transformations, i.e.,
@@
The minimum of Eq. (2) is uniquely determined by solving a Euler-Poincaré differential equation (EPDiff) [3, 39] with a given initial condition of velocity fields, noted as $v_{n}(0)$,
@@
where the operator $D$ denotes a Jacobian matrix, $\operatorname{div}$ is the divergence, and $\cdot$ represents an element-wise matrix multiplication. This is known as the geodesic shooting algorithm [57], which nicely proves that the deformation-based shape descriptor, $\phi_{n}$, can be fully characterized by an initial velocity field, $v_{n}(0)$.
@@
III Our Method: MGAug
This section presents a novel algorithm, MGAug, that learns a multimodal latent distribution of geometric transformations represented by the initial velocity fields $v_{0}$. We first introduce a generative model in latent geometric spaces for DA based on VAE, followed by an approximated variational inference. The augmented data is then used to train various image analysis tasks, i.e., image classification or segmentation. Similar to [65, 6, 7], we employ a joint learning scheme to optimize the MGAug and IAT network parameters. To simplify the notation, we omit the time index of $v_{0}$ in the following sections.
III-A Generative Model in Latent Multimodal Geometric Spaces
Given a group of $N$ input images $\mathbf{J}$ with a template $I$, we assume the latent geometric transformations $\mathbf{x}$ are generated from a number of $C$ clusters. The distribution of $\mathbf{x}$ can be formulated as a Gaussian mixture with means $\boldsymbol{\mu}$ and variances $\mathbf{\Sigma}$, specified by a VAE encoder parameterized by $\beta$, i.e., $p_{\theta}(\mathbf{x}|{\boldsymbol{\epsilon}})$, where $\boldsymbol{\epsilon}$ is a random number generated from $\mathcal{N}(0,\mathbb{I})$. That is to say, the VAE encoder outputs a set of $C$ means $\boldsymbol{\mu}$ and $C$ variances $\mathbf{\Sigma}$. A one-hot vector $\mathbf{z}$ is sampled from the mixing probability $\mathbf{\boldsymbol{\pi}}\in[0,1]$ with $\sum_{c=1}^{C}\mathbf{\boldsymbol{\pi}}^{c}=1$, which chooses one component from the Gaussian mixture. In this paper, we set the parameter $\mathbf{\boldsymbol{\pi}}=C^{-1}$ to make the variable $\mathbf{z}$ uniformly distributed, which is similar to [16]. The initial velocity field $\mathbf{v}$ is generated from a VAE decoder parameterized by $\theta$ and the continuous latent variable $\mathbf{x}$. We parameterize the $\mathbf{v}$ in low-dimensional Fourier space. The entire process to generate the observed image data $\mathbf{J}$ from a set of latent variables $\{\mathbf{v},\mathbf{x},\mathbf{\boldsymbol{\epsilon}},\mathbf{z}\}$ is:
@@ III-A Generative Model in Latent Multimodal Geometric Spaces
Similar to [62, 63], we define the covariance matrix $\boldsymbol{\Sigma}$ as an inverse operator of the differential operator, $L$, which is used in the registration regularity in Eq. (2).
@@
III-B Variational Lower Bound
@@
Using previously defined Eq. (10) and Eq. (11), the posterior $q(\mathbf{v}_{i}|\mathbf{J}_{i},I,\boldsymbol{\lambda})\propto p(\mathbf{J}|\mathbf{v},I,\mathbf{\boldsymbol{\lambda}})\cdot p(\mathbf{v}|\mathbf{x})$ is formulated by Baye’s rule. Following [29], we can derive the posterior $p_{\beta}(\mathbf{z}|\mathbf{x},\mathbf{\boldsymbol{\epsilon}})$ as
@@
$-\mathbf{E}_{q(\mathbf{x}|\mathbf{J})\,q(\boldsymbol{\epsilon}|\mathbf{J})}[KL(p_{\beta}(\mathbf{z}|\mathbf{x},\boldsymbol{\epsilon})\,||\,p(\mathbf{z}))]$ (14)
III-C Joint Learning of MAGug With Image Analysis
We demonstrate the effectiveness of MGAug in image analysis tasks (IAT) by developing a joint learning framework. It has been shown in [65, 6, 7] that these two tasks can be mutually beneficial. The MGAug effectively synthesizes image samples to augment the training dataset, while the IAT can provide extra information on rectifying the learned distributions of latent geometric spaces. An overview of our proposed network architecture is shown in Fig. 2.
Figure 2: An overview of our model MGAug for image analysis tasks.
@@ III-C Joint Learning of MAGug With Image Analysis
MGAug is mainly designed based on the previously introduced inference model, where the approximate posteriors are simulated by neural networks. Analogous to [32], we let the encoder of MGAug, parameterized by $\beta$, approximate the posterior as a mixture of multivariate Gaussians. In the latent space $\mathbf{z}$, we use non-linear hidden layers to learn the parameters (i.e., means and covariances) of the mixture Gaussians. The decoder, parameterized by $\theta$, outputs augmented velocity fields $\mathbf{v}_{n}$, followed by generating their associated deformations $\phi_{n}$ (via Eq. (3)) and augmented deformed images. In this paper, we adopt UNet architecture [50] as the network backbone for MGAug. However, other networks such as UNet++ [66] and TransUNet [8] can be easily applied.
@@
$+\frac{1}{2}(\mathcal{L}\mathbf{v}(\beta,\theta),\mathbf{v}(\beta,\theta))+\text{reg}(\beta,\theta),$
@@
For segmentation, we employ a loss function as a combination of Sørensen--Dice coefficient [15] and cross-entropy (CE). Let $Q$ denote the number of segmentation labels, $S_{q}$ denotes the augmented ground truth, and $R_{q}$ denotes the prediction with the associated probabilistic map $R^{p}_{q}$ for $q$-th label. We formulate the segmentation loss as
@@
The total loss of our joint learning framework of MGAug and IAT is $\ell=\ell_{\text{MGAug}}+\ell_{\text{IAT}}$. We employ an alternative optimization scheme [46, 6] to minimize the loss by alternating between the training of the MGAug network and the IAT network without sharing weights, making it an end-to-end network architecture [65, 7].
IV Experimental Evaluation
@@
IV-A Datasets
2D silhouette shapes. We choose Kimia216 [51], which is a relatively smaller dataset with seven classes of 2D silhouette shape data (bone, bottle, children, face, fish, heart, and key). Each shape class includes approximately $20$ binary images, with the size of $28^{2}$. The total number of our training images for multi-classification is about $140$.
2D handwritten digits. We use $70000$ handwritten digits of $10$ classes (00-$9$) from MNIST repository [35]. Each class includes approximately $7000$ images with the size of $28^{2}$.
3D brain MRIs. We include $803$ subjects of public T1-weighted longitudinal brain MRIs with manually delineated anatomical structures from Open Access Series of Imaging Studies (OASIS) [34], Alzheimer’s Disease Neuroimaging Initiative (ADNI) [27], Autism Brain Imaging Data Exchange (ABIDE) [14], and LONI Probabilistic Brain Atlas Individual Subject Data (LPBA40) [52]. Due to the difficulty of preserving the diffeomorphic property across individual subjects with large age variations, we carefully evaluate images from subjects aged from $60$ to $90$. All MRIs were pre-processed as $256\times 256\times 256$, $1.25mm^{3}$ isotropic voxels, and underwent skull-stripping, intensity normalization, bias field correction and pre-alignment with affine transformations.
IV-B Experiments and Implementation Details
We validate the proposed model, MGAug, on both image classification and segmentation tasks. As MGAug falls into the category of transformation-based DA schemes, we mainly focuses on comparing its performance with similar baselines: randomly generated affine transformations (ATA, a commonly used augmentation strategy), and two state-of-the-art unimodal generative models based on diffeomorphic registrations (AdvEigAug or AEA, an augmentation model based on the statistics of a learned latent space of velocity fields [48]; and UDA, a unimodal deformation-based augmentation using random stationary velocity fields [13]).
Evaluation of classification on small datasets. We first validate our model and other baselines on a relatively small dataset of 7-class 2D silhouette shapes. We adopt six classification network backbones, including a three-layer CNN, ResNet18 [21], VGGNet16 [54], SqueezeNet [26], DenseNet121 [25], and Vision Transformer (ViT) [17].